a business person invests $3,620 at 1% compounded semiannually for two years. What is the future value of the investment, and how much interest will they earn over the two year period?
Answer #1
A B C D E F G
2
3 Amount invested $3,620
4 Interest rate 1%
5 Period 2 Years
6 Compounding Semi-annual
7
8 Semi-annual interest rate 0.50% =D4/2
9 Semi-annual Period 4 =D5*2
10
11 Future Value =$3650*(1+0.50%)4
12 $3,692.94 =D3*((1+D8)^D9)
13
14 Hence the future value is $3,692.94
15
16 Interest earned =Future Value - Amount Invested
17 =$3692.94 - $3620
18 $72.94 =D14-D3
19
20 Hence interest earned is $72.94
21